# The Percept-V Challenge: Can Multimodal LLMs Crack Simple Perception Problems?

### v2

@@ A.1.21 Locate Circles Shape
- This is a grid which contains some coloured circles , rectangles and triangles.
- The rows and columns are numbered in the grid.
- The coordinate of a cell is given by the row number followed by the column number.
- Write down all the coordinates of the cells which contain a green circle.
- The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle.
- The coordinates should be in the format (row,column) with the comma and bracket.
@@ A.1.22 Match Outline
- These are two images of some sequence of shapes separated by black horizontal lines
- The first image consists of a sequence of shapes and the second image consists of a sequence of outlines of shapes
- Determine the number of shapes in the first image that match the outlines in the corresponding place in the second image.
- The output must be given in a single line in the form of ANSWER:x where x is the number of matching shapes
@@ A.1.23 Match Shadow
- These are two images of some sequence of shapes separated by black horizontal lines.
- The first image consists of a sequence of shapes and the second image consists of a sequence of shadows of shapes
- Determine the number of shapes in the first image that match the shadows in the corresponding place in the second image.
- The output must be given in a single line in the form of ANSWER:x where x is the number of matching shadows.
@@ A.1.24 Maze Solving
- This is a maze. - You start from the cell labelled S.
- Your goal is to reach cell labelled E through the shortest path possible.
- You can move up, down, left or right.
- You cannot move through the black walls.
- You can only move through white cells.
- Give the sequence of cells taken to go from S to E.
- Last line of output should be a comma separated line with cell numbers starting with S and ending at E and nothing else.
subsubsectionMirror Image Input Prompt: - These are two images containing different shapes in different colours - The image consists of a vertical blue line at the centre. - Some objects of the first image are mirrored along the given vertical line in the second image. - Other objects in the second image maintain the same postions as in the first image. - Determine the number of objects that are mirrored in the second image. - The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
A.1.25 Numbered Images
@@
- This is an image with some shapes drawn and numbers written on top of the shapes.
- Determine the list of numbers written on top of the circles.
- The output must be given in a single line in the form of CIRCLES:x where x is the list of numbers seen on top of the circles
A.1.26 Sort Circles
@@
- There are several black circles with different sizes.
- Sort these circles by size, from smallest to largest.
- Output must only contain the space separated labels of circles in their sorted order by size.
A.1.27 Sort Lines
@@
- The image has sseveral parallel lines of varying lengths.
- Sort these lines by length, from shortest to longest.
- Last line of the output must only contain the space separated labels of thee lines in their sorted order by length and nothing else.
A.1.28 Vanishing Objects
@@
- These are two images containing shapes such as circles, triangles and squares
- The images are identical except for the fact that the second image has a few objects missing
- Count the number of missing objects in the secnd image
- The output must be given in a single line in the form of COUNT:x
A.1.29 Water Image
@@
- These are two images containing different shapes in different colours
- The image consists of a horizontal blue line at the centre.
- Some objects of the first image are mirrored along the given horizontal line in the second image.
- Other objects in the second image maintain the same postions as in the first image.
- Determine the number of objects that are mirrored along the horizontal line in the second image.
- The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
A.2 Image Generation
@@
A.3 Financial Compensation for Participants

### v3

@@ A.1.21 Locate Circles Shape
- This is a grid which contains some coloured circles , rectangles and triangles. - The rows and columns are numbered in the grid. - The coordinate of a cell is given by the row number followed by the column number. - Write down all the coordinates of the cells which contain a green circle. - The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle. - The coordinates should be in the format (row,column) with the comma and bracket.
Figure 24: Locate Circles Shape
@@ A.1.22 Match Outline
- These are two images of some sequence of shapes separated by black horizontal lines - The two images may also contain a single shape each - The first image consists of rows of shapes and the second image consists of rows of outlines of shapes - Determine the number of shapes in the first image that match the outlines in the corresponding row in the second image. - In case each of the two images contain only a single shape, treat the single shape in each image as a row. - The output must be given in a single line in the form of ANSWER:x where x is the number of matching shapes
@@ A.1.23 Match Shadow
- These are two images of some sequence of shapes separated by black horizontal lines - The two images may also contain a single shape each - The first image consists of rows of shapes and the second image consists of rows of shadows of those shapes. - Determine the number of shapes in the first image that match the shadows in the corresponding row in the second image. - In case each of the two images contain only a single shape, treat the single shape in each image as a row. - The output must be given in a single line in the form of ANSWER:x where x is the number of matching shadows
@@ A.1.24 Maze Solving
- This is a maze. - You start from the cell labelled S. - Your goal is to reach cell labelled E through the shortest path possible. - You can move up, down, left or right. - You cannot move through the black walls. - You can only move through white cells. - Give the sequence of cells taken to go from S to E. - Last line of output should be a comma separated line with cell numbers starting with S and ending at E and nothing else.
Figure 25: Maze Solving
A.1.25 Mirror Image
@@
- These are two images containing different shapes in different colours - The image consists of a vertical blue line at the centre. - Some objects of the first image are mirrored along the given vertical line in the second image. - Other objects in the second image maintain the same postions as in the first image. - Determine the number of objects that are mirrored in the second image. - The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
Figure 26: Mirror Image
Figure 27: Mirror Image
A.1.26 Numbered Images
@@
- This is an image with some shapes drawn and numbers written on top of the shapes - Determine the list of numbers written on top of the circles. - The output must be given in a single line in the form of CIRCLES:x where x is the list of numbers seen on top of the circles
Figure 28: Numbered Images
A.1.27 Sort Circles
@@
- There are several black circles with different sizes. - The image may also have a single black circle. - Sort these circles by size, from smallest to largest. - If there is only a single circles, give the number written on that circle. - Output must only contain the space separated labels of circles in their sorted order by size.
Figure 29: Sort Circles
A.1.28 Sort Lines
@@
- The image has several parallel lines of varying lengths. - The image may also contain only a single line. - Sort these lines by length, from shortest to longest. - If there is only a single line, give the number of that line. - Last line of the output must only contain the space separated labels of thee lines in their sorted order by length and nothing else.
Figure 30: Sort Lines
A.1.29 Vanishing Objects
@@
- These are two images containing shapes such as circles, triangles and squares - The images are identical except for the fact that the second image has a few objects missing - Count the number of missing objects in the second image - The output must be given in a single line in the form of COUNT:x
Figure 31: Vanishing Objects
Figure 32: Vanishing Objects
A.1.30 Water Image
Input Prompt:
- These are two images containing different shapes in different colours - The image consists of a horizontal blue line at the centre. - Some objects of the first image are mirrored along the given horizontal line in the second image. - Other objects in the second image maintain the same postions as in the first image. - Determine the number of objects that are mirrored along the horizontal line in the second image. - The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
Figure 33: Water Image
Figure 34: Water Image
A.2 Overall Performance in Human Study
Model Overall performance
GPT-5-mini 56.81%
GPT-4o 31.94%
o4-mini 56.83%
Gemini 55.58%
Qwen 14.69%
Deepseek 7.08%
Human 95.43%
Table 5: Overall performance in human study
A.3 Image Generation
@@
A.4 Financial Compensation for Participants
@@ A.6 GPU Hours
A.7 One-shot Results
Skill / Domain GPT-5-mini GPT-4o o4-mini Gemini Avg. Acc.
— Performance by Skill —
Visual Discrimination 51.46 11.17 52.51 43.57 39.68
Visual Memory 46.55 20.93 45.96 39.38 38.2
Visual Sequential Memory 38.61 12.33 45.48 33.42 32.46
Visual Figure Ground 63.9 20.0 63.9 48.83 49.16
Visual Form Constancy 53.06 16.64 45.63 40.7 39.01
Visual Closure 40.58 17.75 50.76 40.96 37.51
Visual Spatial Relationship 64.61 24.5 61.59 64.11 53.7
Average of skills 51.25 17.62 52.26 44.42 41.39
— Performance by Domain —
change_colour 30.15 11.5 24.62 21.11 21.84
circle_boxes 26.63 21.5 28.64 24.12 25.22
circle_location 55.78 13.0 48.24 54.27 42.82
circle_right_triangle 100.0 59.0 100.0 96.48 88.87
colours_present 4.02 1.5 5.03 1.01 2.89
comparing_size 17.59 2.5 23.62 36.18 19.97
count_coloured_circles 41.71 16.5 35.68 46.73 35.16
counting_circles 57.79 23.5 48.74 53.27 45.82
counting_locations 35.68 9.0 35.18 28.64 27.12
counting_shapes 60.8 20.2 50.75 56.28 47.01
cross_and_knots 100.0 21.0 100.0 91.46 78.11
graph_counting 33.17 3.0 27.14 31.16 23.62
grid_path 51.76 0.5 46.23 0.5 24.75
identifying_shapes 67.84 41.5 91.46 51.76 63.14
inside_circles 94.97 52.0 92.96 95.48 83.85
layered_colours 31.16 25.0 34.17 24.62 28.74
layered_shapes 21.11 12.5 19.1 19.1 17.95
list_colours 40.7 12.0 40.7 36.18 32.4
list_shapes 91.96 68.5 91.96 82.91 83.83
locate_circles_colour 99.5 1.0 98.49 63.32 65.58
match_outline 58.79 17.5 79.4 71.86 56.89
match_shadow 51.26 16.0 70.35 48.24 46.46
maze_solving 48.24 15.0 31.16 46.23 35.16
mirror_image 30.15 17.5 27.14 20.1 23.72
numbered_shapes 81.91 9.5 77.39 53.77 55.64
sort_circles 34.67 16.5 33.17 25.63 27.49
sort_lines 28.64 19.5 23.62 23.62 23.85
vanishing_objects 34.17 6.0 22.11 22.61 21.22
water_image 34.67 12.5 24.12 20.1 22.85
Full Percept-V Dataset 52.13 18.21 51.01 43.79 41.28
Table 6: Comparison of LLM performance across different skills and domains in one-shot setting. All values are percentages.
Figure 35: The overall accuracy of all models in different skills in one-shot setting.